The table below gives the number of parking tickets received in one semester and the GPA for five randomly selected college students who drive to campus. Using this data, consider the equation of the regression line y^=b0+b1x , for predicting the GPA of a college student who drives to campus based on the number of parking tickets they receive in one semester. Keep in mind, the correlation coefficient may or may not be statistically significant for the data given. Remember, in practice, it would not be appropriate to use the regression line to make a prediction if the correlation coefficient is not statistically significant.

Number of Tickets 2 3 5 6 7
GPA 3.9 3.5 3.3 3.1 1.6

Step 1 of 6 :

Find the estimated slope. Round your answer to three decimal places.
Answer #1

Given the table below no of let trekets xp) 213151617. GPA yil 3-9 | 3.5 3-3 31 8.6 This gives the data of college studentawhere ñ & ý are means of reg na no at periods Here in our problem n=5 estudenta) X = 2 +375 +647 5 1 = 23 = 4.6 . :3.943.5 +3Sini? = 123 bebis64.6-514-6) (3008) 123- 5(4.6) = 624 17:2 b= / b = -0.362709 a = y-bă = bo bos. 3.087 10.3627) (46) [bo = 4:so y = (-0.3627) x +407484 Y = m x + C : : Hence estimated slope = bi E = -0°3827 after three decimal places .. Estimate se (